### v1

# Towards Practical Finite Sample Bounds for Motion Planning in TAMP

Abstract
When using sampling-based motion planners, such as PRMs, in configuration spaces, it is difficult to determine how many samples are required for the PRM to find a solution consistently. This is relevant in Task and Motion Planning (TAMP), where many motion planning problems must be solved in sequence. We attempt to solve this problem by proving an upper bound on the number of samples that are sufficient, with high probability, to find a solution by drawing on prior work in deterministic sampling and sample complexity theory. We also introduce a numerical algorithm to compute a tighter number of samples based on the proof of the sample complexity theorem we apply to derive our bound. Our experiments show that our numerical bounding algorithm is tight within two orders of magnitude on planar planning problems and becomes looser as the problem’s dimensionality increases. When deployed as a heuristic to schedule samples in a TAMP planner, we also observe planning time improvements in planar problems. While our experiments show much work remains to tighten our bounds, the ideas presented in this paper are a step towards a practical sample bound.
Keywords:
Motion and Path Planning, Task Planning, Algorithmic Completeness and Complexity
1 Introduction
Sampling-based motion planners, such as PRMs, perform well in complicated configuration spaces of moderately high dimension. It is difficult to determine how many samples a PRM needs to find a solution consistently. This problem appears when sampling-based motion planners are used as subroutines for Task and Motion Planning (TAMP), where many motion planning problems must be solved in sequence. Existing TAMP solvers address this problem by manually selecting a uniform threshold for the motion planner’s computational effort to find a solution. We seek a more principled method of estimating sample thresholds.
We attempt to solve our problem by proving a bound on the number of random samples for a PRM to find a solution. First, we must choose a set of quantitative features (a problem parametrization) of the motion planning problem on which we derive our bound. Two notable problem parametrizations and sample bounds lay the foundation of our work. Tsao et al. suggest using the widths of passages in the configuration space as a parameterization [31]. They derive the number of deterministic samples required to cover the space at a specified density to yield PRMs guaranteed to have paths in these narrow passages. While their parameterization is simple, TAMP systems generally cannot afford to use deterministic sample sets. If a planner returns to a previously attempted motion planning problem, a denser set of samples must be recomputed from scratch (and previous computation is wasted) [25, 14]. Hsu et al. introduce a random sample bound based on a three-parameter problem characterization called expansiveness, which characterizes the visibility properties of individual points and connected sets in the configuration space [12]. Should we compute samples with his bound, iterative search through three parameters if the planner fails to solve the problem is a difficult task. We wish to have a simple problem parameterization, like Tsao et al., but for random sample sets like Hsu et al. that can be incrementally refined.
This paper presents a sample bound that extends Tsao et al.’s radius PRM guarantee (parameterized by narrow passage width) to random samples. The core of our argument draws a connection to work in sample complexity by Blumer et al. [2] that estimates the probability of finding covering point configurations of a specific density. When Blumer et al.’s result is applied in practice, their bound significantly overestimates the number of required samples resulting in a larger PRM than what can tractably be stored. Rather than using the explicit formula stated in the theorem of Blumer et al., we present a novel numerical algorithm that searches over a tighter expression in their proof [2] to compute a tighter number of samples for a given probability of failure. We also attempt to extend our radius PRM result to the KNN PRM, but our analysis falls short of providing a well-defined bound. However, we include the intermediate analysis here to provide a starting point for a future finite-sample bound.
In our experiments, we evaluate the numerical algorithm’s tightness and utility in a TAMP planner. We find that the numerical bounding algorithm is tightest in planar motion planning problems and loosens as the number of dimensions in the configuration space increases. Consequently, the bound also speeds up planning time for planar TAMP problems but shows limited improvement on higher-dimensional problems. Even so, we believe the theoretical machinery and algorithmic ideas presented in this paper are a useful step towards a practical bound.
2 Background and Problem Formulation
Algorithm 1 TampPlanner
1:TAMP problem $p$=(start, goal, …)
[…]
3:for ($skeleton,progress)\in tp.\text{{SkeletonQueue()}}$() do
4:     $success,paths\leftarrow\text{{True}},[]$
5:     for $(X,X_{free},x_{s},x_{g})$in skeleton do,
6:         $G_{PRM},\delta,\gamma\leftarrow\text{{RestoreProgress}}(X,progress)$triangleright $\delta$ is passage width.
7:         $N_{new}\leftarrow\text{{{ComputeNewSamples}}}(\text{{vol}}(X),\delta,\gamma)$$\triangleright$ $\gamma$ is failure prob.
8:         $G_{PRM}\leftarrow\text{{sPrmConstruction}}(N_{new},X,G_{PRM})$
9:         $path\leftarrow\text{{sPrmQuery}}(s,g,G_{PRM})$
10:         $progress.\text{{update}}(X,s,g,G_{PRM})$
11:         if $path\neq\emptyset$ then
12:              $paths.\text{{append}}(path)$
13:         else
14:              $\delta^{\prime}\leftarrow\text{{FinerResolution}}(\delta)$
15:              $skeleton.\text{{UpdateMpProblem}}(X,s,g,\delta^{\prime})$
16:              $tp.\text{{QueuePush}}(skeleton)$
17:              $success\leftarrow\text{{False}}$
18:              break
19:     if $success$ then:
20:         return $paths$
21:return $\emptyset$
We aim to estimate the computational effort required to solve a constituent motion planning problem in a task skeleton proposed by a task planner. Algorithm 1 sketches out the planning routine we use in our problem setting, which we walk through in this section. We use the PRM definition as Karaman and Frazzoli [15], which is also included in the supplemental material.
The TAMP planner (Alg. 1) uses a priority queue of task plans, each of which proposes a sequence of motion planning problems to be solved by the PRM (lines 2-8). A motion planning problem consists of a configuration space $X\subset\operatorname{\mathbb{R}}^{d}$, a collision-free configuration space $X_{free}\subseteq X$, a start pose $x_{s}\in X$, and a goal pose $x_{g}\in X$, which we will denote by the tuple $(X,X_{free},x_{s},x_{g})$. We assume that $X_{free}$ is Borel measurable and can be sampled uniformly directly, so all our results and experiments will be done using $X_{free}$ directly without much reference to $X$. In practice, the measure of $X_{free}$ can be computed efficiently by Monte Carlo integration [35].
Whenever sampling-based motion planners are used as a subroutine in TAMP, the planner must decide how many samples to use in each problem (line 6). Our main objective is to develop a mathematically well-founded bound on the number of samples required for a PRM to solve the problem (with high probability), and then write an efficient implementation of that bound. Our bound depends on the width $\delta$ of the passages in $X_{free}$ to find paths and the probability of failure $\gamma$ of finding a path even if one exists. Generally, a planning algorithm does not know the widths of the passages of a configuration space in advance. As a result, Alg. 1 first computes samples and PRM assuming wide paths and high failure probabilities on the initial attempt to solve each motion planning subproblem. On successive iterations of the same motion planning problem, the algorithm progressively shrinks the path width and failure probability (and thus increases the number of samples; Alg. 1, line 13). We note that in this implementation of Alg. 1, a skeleton containing an infeasible motion planning problem will always remain in the queue (but may not be given a high priority depending on the task planner). Alternatively, the minimum passage width can be set based on the robot’s control infrastructure’s capability to execute a motion plan in a narrow corridor. Should the planner fail to find a motion plan in a skeleton on an iteration using the minimum width, the planner could remove the skeleton from the queue.
The only piece of notation we require is that $B_{r}^{d}(x)$ represents a $d$-dimensional ball with radius $r$ centered at point $x$. If the location of the ball does not matter (if we are only discussing its volume, etc.), we will abbreviate to $B_{r}^{d}$.
3 Random Covering Bounds for Radius PRMs
We begin by first proving a sample bound for radius PRMs. These PRMs approximate the configuration space with a set of points, where a single point represents a ball region with a radius equivalent to the connection radius. Tsao et al. observed that if a set of samples forms a sufficiently dense covering of the configuration space, the sample set will be a good approximation of the entire space. They then showed that building roadmaps from dense covering sets are guaranteed to find paths in narrow passages of a specific width. We introduce the terminology and results to state precisely what we mean by a narrow passage, a covering set of samples, and the relation between the two for the PRM guarantee. Proofs of all results in this section are included in the supplemental material.
Figure 1: Left: An $\alpha$-net, where the points are interpreted as centers of balls, with radius $\alpha$, that collectively cover the rectangle. Middle: An example of a $\delta$-clear path. Right: A depiction of how a covering as dictated by Lemma 1 yields a roadmap that finds paths in narrow passages.
We start with the definition of narrow passages. If there exists a solution path between $x_{s}$ and $x_{g}$, then we can associate a ‘clearance’ with the path, the extent to which the robot can deviate from the path without hitting obstacles:
Definition 1 (Path Clearance)
Let $(X,X_{free},x_{0},x_{g})$ be a motion planning problem. A continuous path $\sigma:[0,1]\to X$ s.t. $\sigma(0)=x_{0}$ and $\sigma(x_{g})=1$ is $\delta$-clear if $B_{\delta}(\sigma(t))\subset X_{free}$ for all $0\leq t\leq 1$.
We can imagine sliding a ball with radius $\delta$ by its center along the path. Should the ball not hit an obstacle, then the path is $\delta$-clear (Fig. 1). The largest possible $\delta$ characterizes the ‘width’ of the passage through which the path passes.
The mathematical term for a covering set of points is called an $\alpha$-net. A common intuitive definition is first to associate each point with a ball region of radius $\alpha$ centered around the point. Should the union of the $\alpha$-balls subsume the entire configuration space (‘covers’ the configuration space), then we call the set of points an $\alpha$-net (Fig. 1). In our case, we work with an equivalent definition:
Definition 2 ($\alpha$-net)
Let $X\subset\operatorname{\mathbb{R}}^{n}$. A finite subset $N\subset X$ is an $\alpha$-net of $X$ if for every $x\in X$ there exists $x_{n}\in N$ such that $||x-x_{n}||_{2}\leq\alpha$.
We now state a claim proved by Tsao et al. [^1] In Tsao et al. [31], they prove a more complicated version of the claim that involves path optimality., which shows that if a set of points forms an $\alpha$-net, then a (radius) PRM with a suitable connection radius will be guaranteed find paths for paths with clearance $2\alpha$.
Lemma 1 ($\alpha$-nets find paths in $2\alpha$-clear passages)
Let $X_{free}$ be a collision-free configuration space. Suppose that $N\subset X_{free}$ is a finite set of samples that forms an $\alpha$-net of $X_{free}$, used to construct a radius PRM with connection radius $4\alpha$.
For all $x_{s},x_{g}\in X_{free}$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the resulting PRM will return a feasible path from $x_{s}$ to $x_{g}$.
A visual depiction of the proof can be found in Fig. 1c (and the full proof can be found in the supplemental material).
Tsao et al. were interested in deterministic algorithms to construct $\alpha$-nets based on layered point lattices [31]. Because TAMP planners like Alg. 1 may return to previously initiated motion planning efforts, we wish to increment previously built PRMs to save and reuse computational effort as much as possible. Deterministic sample sets are constructed with a fixed net radius $\alpha$ in mind and must be reconstructed from scratch if a finer net is required. This is why we use sample sets obtained from random sampling, which can be naturally incremented when an algorithm revisits a motion planning problem.
To determine the probability that a set of random samples forms an $\alpha$-net, Blumer et al., in their work on probably approximately correct learning (Theorem 2.1 [2]), provide a bound, which we take as a lemma.
Lemma 2 (Number of random samples to form an $\alpha$-net)
Let $X_{free}\subset\operatorname{\mathbb{R}}^{d}$ be a collision-free configuration space, and let $\operatorname{\mathrm{P}}$ be the uniform (Lebesgue) measure on $X_{free}$ and $\gamma\in(0,1)$ be a failure probability.
Suppose $N\subset X_{free}$ is a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (1)
then $N$ is an $\alpha$-net of $X_{free}$ with probability $1-\gamma$.
Putting Lemmas 1 and 2 together, we now have the probabilistic bound for radius-PRMs we were looking for:
Theorem 3.1 (Radius PRM Guarantee)
Let $X_{free}$ be the collision-free configuration space.
Suppose $N\subset X_{free}$ be a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (2)
then for all $x_{s},x_{g}\in X_{free}$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the a radius PRM with connection radius $4\alpha$ will return a feasible path from $x_{s}$ to $x_{g}$ with probability $1-\gamma$.
Before discussing ways to apply this bound to write ComputeNewSamples in Alg. 1, we first want to gain intuition about how path clearance and the dimension of the configuration space influence the number of samples. To do this, a brief asymptotic analysis is in order. Using Stirling’s approximation, the asymptotics of the volume of a ball with radius $\delta$ as $d\to\infty$ is $\Omega\left((\delta/2)^{d}\cdot(2\pi e)^{d/2}\cdot d^{-(d+1)/2}\right)$. If we fix a constant probability $\gamma$ and desired path clearance $\delta$, we obtain the following asymptotic expression:
$n=O\left[\frac{C_{d}}{\delta^{d}}\cdot\left(\log\frac{1}{\gamma}+d\log\frac{C_{d}}{\delta^{d}}\right)\right],\quad C_{d}=\frac{2^{d}d^{(d+1)/2}}{(2\pi e)^{d/2}}.$ (3)
We see that the number of samples varies nearly-linearly in $1/\delta$ and a dimension-dependent constant $C_{d}$, and logarithmically in $1/\gamma$. We do not find this result very surprising – a commonly held belief about PRMs is that one must sample enough to cover the entire space in the worst case, which introduces an exponential sample complexity based on dimension. Our sample complexity argument makes this idea very precise and captures a dependence on path clearance that agrees with prior work on narrow passages [11, 16].
We could use Theorem 3.1 directly by using the right side of the inequality to choose the number of samples for ComputeNewSamples in Alg. 1. In practice, this bound is quite loose. The culprit is the maximization operation in Lemma 2. The maximization’s right-hand argument shows the existence of ‘small’ $\alpha$-nets (up to constant factors) by the probabilistic method. The left-hand argument helps us estimate the probability of failure $\gamma$. The issue is that the constant on the right-hand side is too large, and so $\gamma$ must already be very small (some hand computations suggest less than $10^{-3}$ for moderate net radii) for the left-hand argument to dominate.
To find a tighter number of samples, we must bypass the maximization operation. We summarize Blumer et al.’s proof of Lemma 2 (appears as Theorem 2.1[2]) for insight, which estimates the probability that a sample set $N$ does not form an $\alpha$-net:
[…]
$=$ $\mathbf{P}(\exists x\in X_{free}\mbox{ s.t. }N\cap B_{\alpha}^{d}=\emptyset)$ Defn. $\alpha$-net, Supp. Material (5)
$\leq$ $\left(\sum_{i=1}^{d+1}\binom{2|N|}{i}\right)\cdot 2^{-\operatorname{\mathrm{P}}(B_{\alpha}^{d})|N|/2}$ Blumer et al. Thm. A2.1 (6)
$<$ $\left(\frac{2e|N|}{d}\right)^{d}\cdot 2^{-\operatorname{\mathrm{P}}(B_{\alpha}^{d})|N|/2}$ Blumer et al. Prop. A2.1(iii) (7)
Substituting in the sample inequality in Lemma 2 renders the last line smaller than $\gamma$, for $\gamma$ sufficiently small (Lemma A2.4 [2]).
The third line (Eq. 6) is a much tighter bound on the probability that $N$ fails to be an $\alpha$-net. We write a numerical algorithm that searches for the smallest $n=|N|$ such that the third line is below a given probability $\gamma$ (Alg. 2).
Algorithm 2 NumericalSampleBound
1:$1>\gamma>0$, $p=\operatorname{\mathrm{P}}(B_{\alpha}^{d})$ volume of $d$-dim. $\alpha$-ball by uniform measure on c-space $X$.
2:$N_{l}\leftarrow 1$, $N_{u}\leftarrow 1$
3:while $\left(\sum_{i=1}^{d+1}\binom{2(N_{u}+1)}{i}\right)2^{-p(N_{u}+1)/2}<\left(\sum_{i=1}^{d+1}\binom{2N_{u}}{i}\right)2^{-pN_{u}/2}<\gamma$ do
4:     $N_{u}\leftarrow 2N_{u}$triangleright Find interval that contains desired sample count.
5:while $N_{l}+1<N_{u}$ do
6:     $N_{mid}\leftarrow\left\lfloor(N_{l}+N_{u})/2\right\rfloor$
7:     if $\left(\sum_{i=1}^{d+1}\binom{2(N_{mid}+1)}{i}\right)2^{-p(N_{mid}+1)/2}<\left(\sum_{i=1}^{d+1}\binom{2N_{mid}}{i}\right)2^{-pN_{mid}/2}<\gamma$ then
8:         $N_{u}\leftarrow N_{mid}$triangleright Binary search to find the tightest sample count.
9:     else
10:         $N_{l}\leftarrow N_{mid}$
11:return $N_{u}$
Figure 2: Left: A comparison of the sufficient number of samples required to find an $0.5$-net with probability $1-\gamma$ as computed by Lemma 2 and numerical computation done by Algorithm 2. Most notably, the right-hand side in the maximization in Lemma 2 dominates for nearly all $\gamma$ plotted, since the dotted lines are nearly horizontal. The lines pass each other due to poor numerical precision. Right: A depiction of the search conducted by the numerical algorithm.
This algorithm exploits the fact that probability bound in Eq. 6 exhibits a monotonically increasing and then decreasing behavior as a function of the number of samples. The algorithm uses a doubling scheme to first find a conservative upper bound on the sufficient number of samples on the decreasing side of the global maximum. A binary search narrows down the correct sample count (also on the decreasing side of the global maximum, Fig. 2b). A proof of the algorithm’s correctness may be found in the supplemental material. In practice, we find that using our method is substantially tighter than using the result directly from Lemma 2, and allows us to express the scaling relationship of the number of samples with the failure probability $\gamma$ (Fig. 2b). We do find some looseness in the numerical bound, which is why some of the lines cross each other. This occurs because our implementation experiences some numerical issues when converting the large combinatorial sum in Eq. 6 and small $\gamma$ to log-space for comparison.
The runtime of the algorithm is output-sensitive. Suppose $n$ is the output number of samples. The doubling computation will double $N_{u}$ until $N_{u}\geq n$. Then, the binary search will search across the integers $[N_{u}]$. On every iteration of the doubling and search phases of the algorithm, we sum $d$ binomial coefficient computations. A binomial coefficient computation $\binom{m}{k}$ runs in $O(m)$. Putting everything together, we can express the runtime as the summation $2\sum_{i=0}^{\left\lceil\log n\right\rceil}d2^{i}=O(dn)$.
At this point, Alg. 2 can be used to compute the number of samples in ComputeNewSamples. Experiments to verify the tightness of the bound can be found in Section 5.
4 An Attempt to Extend to KNN PRMs
So far, we have focused on radius PRMs. They are a natural place to start because the connection radius gives us a direct relationship to the clearance of paths we wish our PRMs to find. In practice, radius PRMs can only be used with relatively few samples since the degrees of the vertices tend to grow very quickly as the number of samples increases [15]. $K$-Nearest Neighbor (KNN) PRMs are often used as a space-efficient alternative. This algorithm constructs the roadmap graph by attempting to connect a vertex to its $K$-nearest neighbors in the sample set.
A key assumption in Lemma 1 is that the connection radius is sufficiently large to connect the $\alpha$-net samples to form the roadmap. It is tempting to quantify an ‘effective’ connection radius of a KNN PRM, since then our argument to justify Theorem 3.1 would follow exactly. Ultimately, we find this approach leads to a vacuous bound. However, we include our work here to serve as a good starting point for a finite sample analysis of KNN PRMs.
An effective connection radius of a KNN PRM is a quantity that can be measured after sampling and constructing the KNN graph:
Definition 3 (Connection Radius)
Let $X_{free}$ be a configuration space, and let $N\subset X_{free}$ be a set of samples. Suppose that $G=(V,E)$ is a $K$-nearest neighbor roadmap graph on $N$.
We say $r\in\operatorname{\mathbb{R}}^{+}$ is a connection radius of $G$ if $\forall u,v\in V$ if $\left\lVert u-v\right\rVert\leq r$, then $u$ is at most the $K$th nearest neighbor of $v$, or $v$ is at most the $K$th nearest neighbor of $u$.
There are an infinite number of valid connection radii, and we prefer to find the largest radius we can. It turns out that we can provide a sharp upper bound on all possible valid connection radii:
Lemma 3
Let $X_{free}$ be a collision-free configuration space, and let $N\subset X_{free}$ be a set of samples. Suppose that $G=(V,E)$ is a $K$-nearest neighbor roadmap graph on $N$. Let $r>0$. Then
$r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert,$ (8)
if and only if $r$ is a connection radius of $G$, where $\operatorname{NN}_{i}(v)$ refers to the $i^{th}$ nearest neighbor of vertex $v$.
We also note that Eq. 8 is a conservative approximation for the true connectivity of the KNN PRM (Fig. 3a).
The upper bound in Eq. 8, as a function of a random geometric graph of degree $k$, is a random variable, so we must understand its behavior before we use it as a connection radius as in the analysis for radius PRMs. Fortunately, as the number of samples increases, the realizations of the upper bound tend to concentrate around similar values222 The right-hand quantity in Eq. 8 obeys concentration of measure phenomena, e.g. a generalization of how sums of i.i.d. random variables tend to only deviate from their means by small amounts as stated in the Central Limit Theorem [35]. . We will exploit this fact to provide a conservative lower bound to the expression in Eq. 8 to study how the connection radii of a KNN PRM shrinks as the number of samples increases.
Proposition 1
Let $V$ be a set of $n$ i.i.d. uniform samples from collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM, and suppose that $18K<n-$ 1.333The reason we state this mild technical condition (it is common to sample far more than the number of neighbors) is explained in the appendix. Then
$\left[\frac{K-\sqrt{2K(\log n-\log\gamma)}}{(n-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d}<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert,$ (9)
with probability at least $1-\gamma$.
We now use the lower bound stated in Eq. 9 to stand in for the connection radius assumption in the statement of Lemma 1:
Corollary 1
Let $V$ be a set of $n$ i.i.d. uniform samples from the collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM, and suppose that $18K<n-$ 1. Let $r\in\operatorname{\mathbb{R}}^{+}$. If:
$r\leq\left[\frac{K-\sqrt{2K\log(n/\gamma)}}{(n-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d},$
then $r$ is a connection radius with probability at least $1-\gamma$.
To help parse the bound, we observe the following about the decay rate of the effective connection radius as a function of the failure probability $\gamma$ and number of samples:
  • As the failure probability $\gamma$ increases, the effective connection radius grows on the order of $O\left(\frac{K\log(1/\gamma)}{\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/2d}$.
  • As the number of samples $n$ increases, the effective connection radius decays on the order of $o\left(\frac{K}{n\cdot\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/d}$.
We suspect that the conservative $\sqrt{\log n}$ factor (which makes the bound vacuous as $n\to\infty$ and the culprit for the small $o$ above) can be improved with a tighter analysis. However, for small $n$ (less than 1,000,000), we empirically find that the bound is fairly tight (numerical experimental results are included in the appendix). Furthermore, we can even define a similar numerical algorithm to Alg. 2 that can search for a tighter radius as a function of the number neighbors, samples, and failure probability in the appendix.
Corollary 1 tells us enough about the effective connection radius for us to apply Lemma 1. It is tempting to combine them with a union bound to make an analogous statement to Theorem 3.1 for KNN PRMs:
Theorem 4.1 (Vacuous KNN PRM Guarantee)
Let $X_{free}$ be the collision-free configuration space, and suppose Near is defined be a $K$-nearest neighbor search.
Suppose $N\subset\mathcal{X}$ be a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{4}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (10)
where
$\alpha\leq\frac{1}{4}\cdot\left[\frac{K-\sqrt{2K\log(2n/\gamma)}}{(n-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d},$ (11)
then for all $x_{s},x_{g}\in X$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the PRM will return a feasible path from $x_{s}$ to $x_{g}$ with probability $1-\gamma$.
Figure 3: Left: A depiction of the effective connection radius for a PRM graph with $K=1$. Note that we have a conservative approximation that fails to capture ‘long distance edges’ between isolated vertices. Right: A plot for the ratio of the connection radius to the net radius as a function of the number of sampled vertices with a fixed failure probability of $\gamma=0.1$ and $K=256$ (see Theorem 3.1). The conditions for Theorem 3.1 are satisfied for all samples which have ratios above 4.0 (dotted green line). All of the curves are below the line, so the conditions of Theorem 4.1 are not satisfied for any realistic $n$.
Equation 10 expresses the same sample condition from Lemma 2, and Equation 11 fulfills the connection radius condition stated in the hypothesis of Lemma 1. Unfortunately, as shown by Fig. 3, these two conditions can never be satisfied at the same time. The effective connection radius decays at a rate faster than $o(1/n^{1/d})$, while the largest net radius to be found after $n$ samples is $\Omega(1/n^{1/d})$. Lemma 1 stipulates that these two quantities must form a net-radius to connection-radius ratio of 1:4, which is never true (Fig. 3).
We suspect the vacuousness of the bound comes from two issues: looseness in the analysis of the effective radius and looseness in the union bound combining Prop. 1 and Theorem 3.1. Improvement in both steps of the argument may result in a sample bound for KNN PRMs.
5 Experiments
We seek to experimentally verify the tightness of our sampling bound of convergence to path feasibility by Algorithm 2. Since the conditions on Theorem 4.1 prevent us from formally bounding the number of samples required by a KNN PRM, we explore using the radius bound (Theorem 3.1 and Alg. 2) as a possible sampling heuristic for KNN PRMs, since KNN connection approach is better suited to higher-dimensional problems with many samples.
5.1 Bound Tightness in a Narrow Passageway Problem
Our clearance experiments are conducted in a ‘narrow hallway’ problem that can be generalized to arbitrarily high dimensions. The configuration space can be characterized by
$\underbrace{[-1.5,-0.5]\times[-0.5,0.5]^{d-1}}_{\text{"left end"}}\cup\underbrace{[0.5,1.5]\times[-0.5,0.5]^{d-1}}_{\text{"right end"}}\cup\underbrace{[-0.5,0.5]\times[-\delta,\delta]^{d-1}}_{\text{hallway}}.$
The hallway is the narrow passage and its center path $[-0.5,0.5]\times\{0\}^{d-1}$ has clearance $\delta.$ We construct many instances of this hallway problem for varying $\delta<0.5$ and for $d=\{2,3,6,7,14,20\}$. We form a Monte Carlo estimate of the probability of finding a path in the narrow passages by building 100 PRMs and checking if they return a solution to the query $x_{s}=(-0.5,0,\dots,0),x_{g}=(0.5,0,\dots,0)$ for samples $n=100,1000,10000$. We then compare the number of samples the numerical bound (Alg. 2) predicts to find a path in the passage with the Monte Carlo estimated probability. To explore the use of Alg. 2 as a sampling heuristic for KNN PRMs, we compute the bound according to the clearance of the environment and set the probability of failure to $\gamma=0.99$. The results are in Table 1.
$\delta\downarrow$ Dim ($d$) $n=100$ (rad.) $n=1000$ (rad.) $n=100$ (KNN) $n=1000$ (KNN)
0.499 2 1.00/1.42e+05 -/- 1.00/1.19e+03 -/-
3 1.00/5.13e+06 -/- 1.00/5.20e+03 -/-
4 1.00/2.00e+08 -/- 1.00/2.46e+04 -/-
5 1.00/8.20e+09 -/- 1.00/1.24e+05 -/-
6 1.00/3.53e+11 -/- 1.00/6.60e+05 -/-
0.25 2 1.00/1.17e+05 -/- 1.00/4.53e+03 -/-
3 0.93/3.19e+06 1.00/3.79e+06 1.00/3.73e+04 -/-
4 0.37/1.21e+08 1.00/1.39e+08 1.00/3.45e+05 -/-
5 0.03/5.02e+09 1.00/5.57e+09 1.00/3.45e+06 -/-
6 0.00/2.19e+11 0.83/2.21e+11 0.95/3.67e+07 0.96/3.67e+07
0.125 2 0.19/7.37e+04 1.00/1.05e+05 1.00/1.86e+04 -/-
3 0.00/2.78e+06 0.95/2.93e+06 0.58/3.24e+05 0.99/3.24e+05
4 0.00/1.14e+08 0.00/1.14e+08 0.09/6.36e+06 0.60/6.36e+06
5 0.00/4.87e+09 0.00/4.87e+09 0.02/1.33e+08 0.02/1.33e+08
6 0.00/2.16e+11 0.00/2.16e+11 0.00/2.89e+09 0.00/2.89e+09
0.0625 2 0.00/6.87e+04 0.95/7.53e+04 0.86/7.88e+04 1.00/7.88e+04
3 0.00/2.71e+06 0.00/2.71e+06 0.03/2.93e+06 0.53/2.93e+06
4 0.00/1.13e+08 0.00/1.13e+08 0.00/1.19e+08 0.00/1.19e+08
5 0.00/4.86e+09 0.00/4.86e+09 0.00/5.04e+09 0.00/5.04e+09
6 0.00/2.15e+11 0.00/2.15e+11 0.00/2.21e+11 0.00/2.21e+11
[…]
For radius PRMs, we find that the bound is loose 2-3 orders of magnitude for the planar hallway problem and becomes looser as the number of dimensions increases. In KNN PRMs, the radius bound (as a sampling heuristic) is tight for 1-2 orders of magnitude and loosens as the number of dimensions grows. This means other conditions besides the sufficient conditions outlined in Lemma 1 can allow the PRM to find a path through the narrow passage. Surprisingly, we find that Alg. 2 is better suited as a sampling heuristic in lower dimensions for KNN PRMs than for radius PRMs.
5.2 TAMP Experiments
Our experiments in Table 1 suggest that the radius PRM bound may be a useful sampling heuristic for a KNN PRM, which is more tractable to compute in practice. We now aim to assess the usefulness of such a bound in the context of a TAMP planner that must solve multiple motion planning subproblems.
(a) Our 2DoF mobile-base manipulation problem setting (left) and a scatter plot showing TAMP success vs planning time (right) for the three PRM-based motion planning methods. Each point on the scatter plot is the performance and planning time across 100 environment randomizations in which we vary the number and size of objects. The iPRM and sPRM points represent different settings of N.
(b) Our 6DoF peg-in-hole problem setting (left) and corresponding baseline evaluations (right).
We examine the bound’s performance in a planar task involving a circular robot with 2DoF and its scalability to higher dimensions as we experiment with a 6DoF peg-in-hole insertion task. In the task shown in Figure 4(a), a mobile-base robot must transfer all yellow objects (5-10) of varying sizes from one room to another by passing through a narrow passage. All objects are made to be small enough to fit through the hallway. In Figure 4(b), a 6DoF KUKA robot arm must move all objects (5-10) of varying size from an initial region into a packing container on the right. Each robot’s planner must solve a sequence of motion planning subproblems of varying difficulty based on the size of the object the robot is holding. We make use of the Focused PDDLStream TAMP algorithm [9], which proposes high-level mode sequences (e.g., object grasps and releases) and calls a motion planning subprocess to refine the task skeleton or deem it infeasible.
Using an adaptive bound selection strategy, which we will call aPRM, we compute an upper bound for the number of samples PRM uses upon request from the TAMP solver. Given the size of the object being held, we are able to compute $\delta$ by subtracting the hallway size from the extent of the robot when holding the object, which is sometimes larger than the extent of the robot itself. We then call sPRM outlined by Karaman and Frazzoli [15] (see appendix) with an $N$ determined by the numerical bound for radius PRMs (Alg. 2).
We compare this strategy to two other baseline strategies. First, we compare to the sPRM motion planner with a fixed N selected by hand and fixed across subproblems, which is standard practice in TAMP. Second, we compare to an incremental version of sPRM, which we call iPRM. iPRM iteratively calls sPRM with an increasing max sample bound $K$, which increases geometrically according to $M_{i+1}=cM_{i}$ until a motion plan is found or $M\geq{N}$. Our experiments use $c=1.1$. We compare both of these strategies with aPRM in Figure 4(a).
We find that in the planar problem, we observe that aPRM is able to solve nearly all problems with less computation time than sPRM and iPRM. The radius numerical bound helps the planner effectively trade sampling effort with the complexity of the motion planning problems, whereas sPRM either undersamples or oversamples problems with a single threshold, and iPRM wastes computation time building up to the correct number of samples. In the 6DoF problem, our results are weaker, where some instances of iPRM scale up to a sufficient number of samples faster than aPRM takes in a single computation. This is consistent with our observation that the numerical bound becomes looser as the number of dimensions increases.
6 Related work
Another approach to terminating sampling-based motion planners is to decide the motion planning problem’s feasibility directly. In the literature, there are two common classes of methods to solve this problem. One common approach is to show the existence of a ‘disconnection proof,’ an obstacle region that fully encloses the goal from an outer starting configuration [1, 18, 19]. These proofs usually consist of discrete representations of hypersurfaces that fully enclose the region around the goal. Others used volumetric approximations of the configuration space or obstacle region [33, 23, 36] to find a path or refute feasibility (up to the accuracy of the approximation). These methods scale poorly with a dimension of the configuration space and often empirically are intractable in larger than four dimensions. These representations are computationally expensive to construct, and in most situations, evaluating a cheaper probabilistic bound may be more desirable.
Much of our understanding of PRMs with random sampling is from asymptotic results stated in the limit of infinite samples. Karaman and Frazzoli [15] proved that paths returned by PRM (and the more space-frugal PRM*) are both asymptotically optimal. Tighter estimates on the necessary connection radius to ensure graph connectivity have been obtained as well [28]. All of these results are proven using tail-field type arguments that are difficult to extend to a finite number of samples.
In addition to Hsu et al.’s work on configuration space expansiveness [11], there are several instances of finite-random sample bounds. Kavraki et al. [16] and Dobson et al. [7] gave a bound on the number of samples required to return a path and converge to a near-optimal path, respectively, by assuming and clearance and path length. They both make stronger assumptions on the solution, like path length, for simpler arguments that do not rely on covering sets. These assumptions also preclude them from application in scenarios where the knowledge of the solution path length is unknown. Deterministic low-discrepancy and low-dispersion sampling processes are sample sets that cover the space with a prespecified density. They are justified by superior bounds on sample dispersion (largest unsampled ball) as number of samples increases [25] and present better empirical performance in PRMs than random samples [14]. However, their performance gains over random samples degrades as dimensions increase [13], and while incremental sequences exist, the best-performing sequences are non-incremental [25, 21]. There are techniques that layer varying-density deterministic sample sets to produce an incremental sequence [20], but at the cost of repeating computational effort from coarser grids. For these reasons, random samples are a suitable choice for our problem setting.
Outside of motion planning theory, there are sample bounds for random coverings have been studied in discrete geometry and computational learning theory [22]. The techniques to prove bounds of these kinds has been credited to Vapnik and Chernonenkis [32], Sauer [26], and Shelah [27], who sought to find small coverings by the probabilistic method [^4] Please see the appendix for more details of how these details fit into the proofs of our results.. In computational learning theory, Blumer et al. estimate the probability that the samples form a covering [2] to form probably approximate correct learning guarantees. Their results for the smallest $\alpha$-net (right-hand argument in Lemma 2) were confirmed to be tight up to constants by Kómlos et al. [17] by a matching lower bound. This paper draws a connection between these results and motion planning.
Lastly, this paper shows the potential usefulness of a sampling bound in a search-then-sample style planner. Search-then-sample is a fairly common approach to TAMP [9, 5, 6]. Such planners use sampling-based motion planners use ad hoc sample thresholds or runtimes, which we show can be replaced with a sample bound. In addition to setting resource limits on individual motion planning subproblems, a sample bound could be used to guide the task-level search as a heuristic to decide which motion planning problems to attempt. Some related works along these lines include Cambon et al., wherein the number of failed queries is a heuristic to decide which task plans to prioritize for motion computation [4] and Srivastava et al., where estimated subproblem feasibility under different object pose configurations is used to discover obstacles that must be moved to construct a multi-step plan [29]. We do note that some methods do not directly benefit from a sample bound by the nature of their construction. Some TAMP planners explore by constructing motions first [8, 34], and only make task decisions based on constructed motion plans. Others forgo sampling-based motion planning completely in favor of optimization-based obstacle avoidance [30].
7 Conclusion
We present a theoretical bound for the number of samples sufficient for a radius PRM to find a path of a specified clearance with high probability. Our results are derived by connecting prior analysis of the radius PRM to results from sample complexity theory. Since the results from sample complexity theory significantly overestimate the number of samples required to find a solution, we introduce a novel algorithm that computes a tighter number of samples. We also made progress in extending our theoretical results to the KNN PRM by studying the decay rate of the ‘effective connection radius’ as the number of samples increases.
Our experiments show that despite being tighter than the original theoretical bound, our numerical sample bound is still too loose for practical use. In settings with a single narrow passage, we saw that the bound is tight within two orders of magnitude for low dimensions ($d=2$) and becomes increasingly loose as the number of dimensions grows. We then applied our numerical bound (for radius PRMs) as a heuristic to a TAMP planner equipped with a KNN PRM in problems where the narrow passage width was known. Our results show that only in low dimensional problems ($d=2$), the bound was a useful heuristic to help the planners scale the number of samples appropriately, given the complexity of the motion planning problem.
In the future, we wish to tighten our numerical bound for more practical use in higher dimensions. We can further explore Blumer et al.’s proof for improvements: our numerical algorithm only tightens Blumer’s bound over a single lossy step, and there are many more such steps that may lead to a significantly tighter bound. We also plan to finish extending the sample bound to KNN PRMs to bridge the gap between our theory and experiments, by investigating the connection between sampling to find $\alpha$-nets and effective connection radii.
7.0.1 Acknowledgements
We thank David Hsu, Sandeep Silwal, Justin Chen, Haike Xu, Josh Engels, Thomas Cohn, and Christopher Bradley for their invaluable support and feedback.
Supplemental Proofs, Algorithms, and Experiments
A.1 PRM Pseudocode
Algorithm 3 sPrmConstruction
1:$n>0$ number of new samples, $X$ c-space, $G=(V,E)$ prev. PRM graph
2:$V$ $\leftarrow V\cup\{\textsc{SampleFree}(X)\}_{i=|V|,\dots,n}$
3:for $v\in V$ do
4:     $U\leftarrow\textsc{Near}(v,V\backslash\{v\})$
5:     for $u\in U$ do
6:         if $\textsc{CollisionFree}(v,u,X)$ then
7:              $E\leftarrow E\cup\{(u,v)\}$
8:return $(V,E)$
Algorithm 4 sPrmQuery
1:$x_{s}$ start, $x_{g}$ goal, $X$ c-space, $G=(V,E)$ PRM graph
2:$U_{s},U_{g}\leftarrow\text{{Near}}(s,V),\text{{Near}}(g,V)$
3:for $u\in U_{s}$ do
4:     if $\textsc{CollisionFree}(x_{s},u,X)$ then
5:         $E\leftarrow E\cup\{(x_{s},u)\}$
6:for $u\in U_{g}$ do
7:     if $\textsc{CollisionFree}(x_{g},u,X)$ then
8:         $E\leftarrow E\cup\{(x_{g},u)\}$
9:return $\text{{ShortestPath}}(G,x_{s},x_{g})$
Karaman and Frazzoli’s [15] implementation of the PRM algorithm uses two separate routines: a pre-processing routine that builds or grows a graph to a specified number of samples (Alg. 3), and a query routine that connects the starts and goals to the graph and runs a shortest-path search (Dijktra, etc., Alg. 4). For both algorithms, we assume several subroutines have already been provided:
  • SampleFree($\cdot$), which samples a collision-free point from the configuration space, and
  • CollisionFree($\cdot$), which checks if a linear path between the query vertices is collision-free.
Two implementations of the Near subroutine can be used: a range search that identifies the vertices within a radius of the query vertex or a $K$-nearest neighbor search.
A.2 Proof of Lemmas 1 and 2 in Section 3
We begin with our proof of Lemma 1, which proceeds along similar reasoning to Tsao et al. [31] (though simpler since we do not discuss path optimality).
Proof (Lemma 1)
Let $x_{s},x_{g}\in X_{free}$ be a start and goal pose, respectively. Suppose there exists a continuous $2\alpha$-clear path $\sigma:[0,1]\to X_{free}$ such that $\sigma(0)=x_{s}$ and $\sigma(1)=x_{g}$.
Let $L$ denote the length of $\sigma$. Let $\{B_{2\alpha}^{d}(p_{i})\}_{i\in[\left\lceil L/{2\alpha}\right\rceil]}$ denote a set of balls on $\sigma([0,1])$, where $B_{2\alpha,1}^{d}(p_{1})$ is centered on $x_{s}=p_{1}$, $B_{2\alpha}^{d}(p_{\left\lceil L/2\alpha\right\rceil})$ is centered on $x_{g}=p_{\left\lceil L/2\alpha\right\rceil}$. The remaining balls (for $i=2,\dots,\left\lceil L/{2\alpha}\right\rceil-1$) are centered along $\sigma([0,1])$, where the $p_{i}$ is spaced $2\alpha(i-1)$ away from $x_{s}$ along arclength of $\sigma$.
Next, we construct another set of balls, $\{B^{d}_{\alpha}(q_{j})\}_{j\in[\left\lceil L/2\alpha\right\rceil-1}]$, where $q_{j}$ is spaced $\alpha+2\alpha(j-1)$ along the arclength of $\sigma$ from $x_{s}=p_{1}$. We observe that there $q_{j}$ is spaced exactly between $p_{j}$ and $p_{j+1}$ along $\sigma$’s arclength.
Since these points are in a Euclidean space, we know that
$\left\lVert p_{j}-q_{j}\right\rVert,\left\lVert p_{j+1}-q_{j}\right\rVert\leq 2\alpha$
for all $j\in[\left\lceil L/2\alpha\right\rceil-1]$. Thus, by construction, we know that
$B^{d}_{\alpha}(q_{j})\subset B^{d}_{2\alpha}(p_{j}),B^{d}_{2\alpha}(p_{j+1}).$
Since $N$ forms an $\alpha$-net, we know that there must exist a point $x_{j}\in N\in B^{d}_{\alpha}(q_{j})$ for all $j$ ($x_{j}$ must be at most $\alpha$ from ball center $q_{j}$). Furthermore, we then know that the linear path between $x_{j}$ and $x_{j+1}$ is collision free for all $j$, since $x_{j},x_{j+1}\in B^{d}_{2\alpha}(p_{j+1})$ and balls are convex.
The maximum distance between $x_{j},x_{j+1}$ is $4\alpha$ (opposite sides of the $2\alpha$-ball, and so setting the connection radius to $4\alpha$ ensures all $x_{s},x_{1},\dots,x_{\left\lceil L/2\alpha\right\rceil-1},x_{g}$ will be connected to form a path to be returned by the radius PRM (the endpoints trivially included as the centers of the first and last $2\alpha$ balls). $\blacksquare$
We now describe the our proof for Lemma 2, Our arguments are centered around covering and point configurations and the probability they arise from random sampling processes, a topic well-studied at the intersection of statistics, combinatorics, and discrete geometry [2, 17, 22, 24] [^5] We will be using nomenclature from both computational geometry and statistics to avoid conflicting terminology..
Our intuition behind the definitions is grounded in trying to ‘hit’ a class of subsets of some space $X_{free}$ with a random finite set of points.
Definition 4
$Y$ be a (potentially infinite) set, and let $\operatorname{\mathcal{F}}\subset 2^{Y}$. We call the tuple $(Y,\operatorname{\mathcal{F}})$ a range space. [^6] For those who are familiar with learning theory: another name for $\operatorname{\mathcal{F}}$ can be the hypothesis space of a set of binary classifiers or set of indicators associated with all elements of $\operatorname{\mathcal{F}}$.
For us, we will take the base space to be $X_{free}\subset\operatorname{\mathbb{R}}^{n}$, the obstacle-free configuration space of a robot. $\operatorname{\mathcal{F}}$ will be the set of $n$-dimensional balls that are a proper subset of $X_{free}$.
Often, hitting all the sets in $\operatorname{\mathcal{F}}$ with a finite number of points in $Y$ is too tall an order to ask. For our purposes, only hitting the ‘voluminous’ sets in $\operatorname{\mathcal{F}}$ is enough.
Definition 5
Let $(Y,\operatorname{\mathcal{F}})$ is a range space. Suppose that $\mu$ is a measure on $Y$ such that every $S\in\mathcal{F}$ is measurable with respect to $\mu$.
A finite subset $N\subset Y$ is an $\epsilon$-transversal if for all $S\in\mathcal{F}$ such that $\mu(S)\geq\epsilon\cdot\mu(Y)$, then $N\cap S\neq\emptyset$.
Intuitively, $\epsilon$ is a threshold on the minimum volume of $S\in\operatorname{\mathcal{F}}$ to be hit by a point in $N$. There could be many different $\epsilon$-transversals on $(Y,\operatorname{\mathcal{F}})$, and we prefer to find smaller $\epsilon$-transversals if we can.
Blumer et al. [2], generalizing Welzl and Haussler’s work [10] from discrete to continuous range spaces showed that random sampling can find $\epsilon$-transversals with high probability.
Theorem A.1
Let $(Y,\mathcal{F})$ be a range space with VC-dimension $\operatorname{VC}(\operatorname{\mathcal{F}})$ and $Y\subset\operatorname{\mathbb{R}}^{n}$. Let $\operatorname{\mathrm{P}}$ be a probability measure on $Y$ such that all $S\in\mathcal{F}$ is measurable with respect to $\operatorname{\mathrm{P}}$.
Let $N\subset Y$ be a set of $m$ independent random samples of $Y$ with respect to $\operatorname{\mathrm{P}}$, where
$m\geq\max\left\{\frac{4}{\epsilon}\log\frac{2}{\gamma},\frac{8d}{\epsilon}\log\frac{13}{\epsilon}\right\},$
then $N$ is an $\epsilon$-transversal of $Y$ with probability at least $1-\gamma$.
Intuitively, the VC-dimension is a quantification of the complexity, or intricacy, of the ways the sets in $\operatorname{\mathcal{F}}$ intersect with $Y$ and each other. We now move to the technical definition:
Definition 6
Let $(Y,\operatorname{\mathcal{F}})$ a range space. We say that a subset $A\subset Y$ is shattered by $\operatorname{\mathcal{F}}$ if each of the subsets of $A$ can be obtained as the intersection of some $S\in\operatorname{\mathcal{F}}$ with $A$. We define the VC-dimension of $\operatorname{\mathcal{F}}$ as the supremum of the sizes of all finite shattered subsets of $Y$. If arbitrary large subsets can be shattered, the VC-dimension is $\infty$.
Our arguments only require the fact that the VC-dimension of a set of $d$-dimensional (Euclidean) balls in $\operatorname{\mathbb{R}}^{d}$ is $d+1$ [22, 24]. When uniformly sampling from a space $X_{free}$, we are in a favorable scenario where $\epsilon$-transversal (hitting set of balls in $X_{free}$) $\epsilon$-net (covering $X_{free}$ with balls) defined in Def. 2 coincide.
Lemma 4
Let $Y\subseteq\operatorname{\mathbb{R}}^{d}$, and $\operatorname{\mathrm{P}}$ be a uniform probability measure. Let $N\subset Y$ be a finite subset.
$N$ is an $\epsilon$-net of $Y$ if and only if $N$ is an $\epsilon^{d}\operatorname{\mathrm{P}}(B^{d}_{1})$-transversal on the range space $(Y,\mathcal{F})$ where $\mathcal{F}$ is the set of $d$-dimensional balls in $\mathbb{R}^{d}$ that have non-empty intersection with $Y$.
Proof
The essence of this proof is a radius-volume conversion, which is only possible since we are working with range spaces with spheres.
We start with the forward direction. Suppose $N$ is an $\epsilon$-net. Let $S\in\mathcal{F}$ such that $\operatorname{\mathrm{P}}(S)\geq\epsilon^{d}\operatorname{\mathrm{P}}(B_{1}^{d})\cdot\operatorname{\mathrm{P}}(Y)=\epsilon^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$ (i.e., $S$’s volume is larger than that of a $d$-dimensional ball with radius $\epsilon$). Since $S$ is a ball, its radius must be larger than $\epsilon$. Let $x_{S}$ be its center. By the definition of geometric $\epsilon$-net, there exists $x_{n}\in N$ such that $||x_{c}-x_{n}||\leq\epsilon$, so $x_{n}\in S$.
The backward direction is even shorter. Let $N$ be a $\epsilon^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$-transversal. Let $x\in\mathcal{Y}$. Let $S$ be the $d$-dimensional ball centered at $x$. Then there must exist $x_{n}\in N$ such that $x_{n}\in S$, so $||x-x_{n}||\leq\epsilon$. $\blacksquare$
We now have all the ingredients we require for Lemma 2:
Proof (Lemma 2)
By Lemma 4 $\epsilon^{d}\operatorname{\mathrm{P}}(B^{d}(1))$-transversals is an $\epsilon$-net. We apply Lemma A.1, and the result follows. $\blacksquare$
A.3 Proofs of Lemma 3 and Prop. 1 in Section 4
The proof of Lemma 3 is a simple proof of contradiction. The intuition for the proof is pictorially displayed in Fig. 3.
Proof (Lemma 3)
We start with the first statement. Let $r\in\operatorname{\mathbb{R}}$ such that $0<r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert$. Suppose, for the sake of contradiction, that there exists $u,v\in V$ such that CollisionFree$(u,v)$ returns True but $(u,v)\notin E$.
Then $v$ must not be a $K$th nearest neighbor of $u$, or vice versa. Without loss of generality, assume the former, and so $v$ must be at least a $(K+1)$th nearest neighbor of $u$. But then $r<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert<\left\lVert u-v\right\rVert$, so we have a contradiction.
The proof of the second statement is fast: suppose $r\geq\min_{v\in V}||v-\operatorname{NN}_{k+1}(v)||$. Let $v^{*}$ be the minimizing vertex in the expression above. Then $(v^{*},\operatorname{NN}_{k+1}(v))$ is a non-edge and $\left\lVert v^{*}-\operatorname{NN}_{k+1}(v^{*})\right\rVert\leq r$. $\blacksquare$
This proof is loosely inspired by the KNN proof for Lemma 58 done by Karaman and Fazzoli in [15], which uses a Chernoff bound to control the number of samples that hit a ball with a radius that scales with the volume and dimension of the environment and number of samples taken:
Proof (Proposition 1)
We prove our bound by controlling the number of samples land within a ball of radius $r$ of another sample using a Chernoff bound.
Let $r=\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert$. Let $I_{v,w}$ be a Bernoulli random variable where vertex $w$ falls within radius $r$ for vertex $v$. We aim to find a lower bound to the probability the event:
$\operatorname{\mathrm{P}}\left(\forall v\in V,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)<K+1\right)$
It is far easier to reason about an upper bound of the complement of this event and take a union bound over all vertices:
$\operatorname{\mathrm{P}}\left(\exists v\in V,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)\geq K+1\right)$
$\leq$ $\sum_{v\in V}\operatorname{\mathrm{P}}\left(\sum_{w\in V\backslash\{v\}}I_{v,w}(r)\geq K+1\right)$
$\leq$ $N\cdot\operatorname{\mathrm{P}}\left(v\in V\text{ s.t. }B_{r}^{d}(v)\subset X,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)\geq K+1\right)$
The second line takes a union bound over all vertices. The third line applies the bound $\operatorname{\mathrm{P}}(B_{r}^{d}(v)\cap X)\leq\operatorname{\mathrm{P}}(B_{r}^{d}(v))$. To obtain a tail estimate on the bimonial random variable above, we use the Chernoff bound stated in [3]:
$\operatorname{\mathrm{P}}\left(v\in V\text{ s.t. }B_{r}^{d}(v)\subset X,\sum_{w\in V\backslash\{v\}}I_{v,w}(r)\geq K+1\right)$
$\leq$ $\exp\left\{-D_{KL}(p+t||p)(N-1)\right\}$
where $D_{KL}$ is the KL-divergence between two Bernoulli variables, and $p=\operatorname{\mathrm{P}}(B_{r}^{d})$ and $t=K/(n-1)-p$ (our bound on $r$ will avoid a degenerate $t\leq 0$).
Using the well-known inequality $D_{KL}(p_{1}||p_{2})\geq(p_{1}-p_{2})^{2}/2p$ when $p_{1}<\frac{1}{8}$ and $p_{1}>p_{2}$ (the mild technical condition stated in the hypothesis) we have that:
$\exp\left\{-D_{KL}(p+t||p)(N-1)\right\}$ $\leq\exp\left\{\frac{-(N-1)t^{2}}{2(p+t)}\right\}$
$\leq\exp\left\{-(N-1)\frac{\left(\frac{K}{N-1}-p\right)^{2}}{2\left(\frac{K}{N-1}\right)}\right\}$
$=\exp\left\{-\frac{(K-p(N-1))^{2}}{2K}\right\}$
Substituting the bound above back into our union bound, plugging in $p=\operatorname{\mathrm{P}}(B_{r}^{d})$, imposing that the last expression be less than $\gamma>0$, and solving for $r$ yields:.
$\operatorname{\mathrm{P}}(B_{r}^{d})=p\leq\frac{K-\sqrt{2K(\log N-\log\gamma)}}{N-1}$
$r\leq\left[\frac{K-\sqrt{2K(\log N-\log\gamma)}}{(N-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d}$ (12)
What we have shown is that if we choose a ball radius $r$ as expressed in Equation 12, then each ball of radius $r$ centered at each vertex in the PRM contains less than $K+1$ other vertices with probability at least $1-\gamma$. By the construction of the event, we know that:
$\left[\frac{K-\sqrt{2K(\log N-\log\gamma)}}{(N-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d}<\min_{v\in V}\left\lVert v-\operatorname{NN}_{k+1}(v)\right\rVert,$
with probability $1-\gamma$. $\blacksquare$
A.4 Numerical Algorithm Results
A.4.1 Analysis of Correctness of Algorithm 2
We begin by stating the following theorem that formalizes the algorithm’s correctness and sketching out the proof:
Theorem A.2
Let $f:\operatorname{\mathbb{Z}}^{+}\to\operatorname{\mathbb{R}}^{+}$ denote the function represents the Eq. 6.
$f(m)=\left(\sum_{i=1}^{d^{\prime}}\binom{2m}{i}\right)\cdot 2^{-cm}$
for $c>0$, and $d^{\prime}\in\operatorname{\mathbb{Z}}^{+}$, which are computed by the input to Alg. 2.
Let the input failure probability be $\gamma\in(0,1)$. Then Alg. 2 will return $m^{*}$ such that $f(m^{*}-1)>\gamma>f(m^{*})$.
Proof
By the definition of the binomial coefficient operation, we know that the left-hand term is a $d^{\prime}$-degree polynomial that is monotonically increasing in $m$ with all positive coefficients:
$\sum_{i=1}^{d^{\prime}}\binom{2m}{i}=\sum_{i=0}^{d^{\prime}}a_{i}m^{i},\mbox{ }a_{i}\geq 0\forall i=0,\dots,d^{\prime}$
We then compute the derivative of $f$. Using the product rule, we see that:
$f^{\prime}(m)=\left(\sum_{i=1}^{d^{\prime}}ia_{i-1}m^{i-1}\right)\cdot 2^{-cm}-cm\left(\sum_{i=0}^{d^{\prime}}a_{i}m^{i}\right)\cdot(2^{-cm}\log 2)$
We observe from the form of the expression (lower order polynomial and exponential on the left term, higher order polynomial and linear factor on the right term) and conclude two possibilities: (1) $f(m)<0$ for all $m\geq 1$ or (2) $f(m)\geq 0$ for some $k>m$, and then $f(m)<0$ for all $m>k$. If we are in the first case, we then write $k=0$ for ease of the rest of the analysis.
In either case, the doubling search of the numerical algorithm will double $N_{u}$ until $N_{u}>k$ and $f(N_{u})\leq\gamma$. Thus, we observe that the binary search will preserve two invariants: (1) $N_{l}\leq m^{*}\leq N_{u}$, since the search increases $N_{l}$ if either (1) $f(\left\lfloor(N_{l}+N_{u})/2\right\rfloor)>\gamma$ or (2) $N_{l}<k$ (by the decrease check on line 6), and (2) the intervals close by half each time.
Thus, since we are searching over a discrete set while the invariants hold, then the algorithm will return the correct answer.
A.4.2 Numerical Algorithm for Proposition 1 and Experimental Results
As mentioned in Section 4, we write an analogous algorithm (Alg. 5) to compute the largest effective connection radius given a failure of probability $\gamma$ (and other technical details of the configuration space). The algorithm performs a binary search up to a prespecified additive approximation error. The correctness of the algorithm follows from a similar argument provided for Alg. 2 above.
Algorithm 5 NumericalRadiusBound
1:$1>\gamma>0$, $\operatorname{\mathrm{P}}(B_{1}^{d})$ volume of unit ball by uniform measure on c-space $X$, $K$ number of neighbors, $n$ number of samples, $r_{u}$ maximum possible radius in $X$, and maximum $\epsilon>0$ additive error of solution.
2:$r_{l}\leftarrow\epsilon$
3:$p\leftarrow K/(n-1)$
4:while $r_{u}-r_{l}>\epsilon$ do
5:     $r_{mid}\leftarrow(r_{l}+r_{u})/2$
6:     $q_{mid}\leftarrow r_{mid}^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$
7:     $q_{mid-}\leftarrow(r_{mid}-\epsilon/2)^{d}\operatorname{\mathrm{P}}(B_{1}^{d})$triangleright Ensure probability bound is decreasing.
8:     if $e^{-D_{KL}(p||q_{mid-})}\geq\exp^{-D_{KL}(p||q_{mid})}$ or $(n-1)\exp^{-D_{KL}(p||q_{mid})}>\gamma$ then
9:         $r_{u}\leftarrow r_{mid}$triangleright Binary search to ensure radius corresponding to $\gamma$ is in $[r_{l},r_{u}]$.
10:     else
11:         $r_{l}\leftarrow r_{mid}$
12:return $r_{u}$
To verify the tightness of Prop. 1 and its numerical algorithm 5, we run an approximate $K$-nearest neighbors algorithm and compute the effective radius upper bound for varying numbers of neighbors, samples, and dimensions. We see that the bound often starts of conservative, and then progressively gets tighter as samples increase (with maximum number of samples of $n=10^{6}$. The rate at which the bound tightens in dependent on the dimension. The rate for $d=2,3$ is quite fast, while the does not fully connverge to tightly bound the right tail of the empirical distributions when $d=20$. Eventually, as $n\to\infty$, we expect the bound will become loose again.
Figure 5: The largest connection radius computed by the numerical bound by setting the probability of failure $\gamma=0.01$ is represent by the blue bars. Their right-to-left ordering is the same as the ordering of the histograms.